# Nonequispaced fast Fourier transforms for bandlimited functions

### v1

@@ 4 NFFT-like procedure for bandlimited functions
$\boldsymbol{D}_{\hat{\psi}}\coloneqq\text{diag}\left(\frac{1}{|\mathcal{I}_{\boldsymbol{L}}|\cdot\hat{\psi}(\boldsymbol{k})}\right)_{\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}}\ \in\mathbb{C}^{|\mathcal{I}_{\boldsymbol{M}}|\times|\mathcal{I}_{\boldsymbol{M}}|},$ (4.7)
@@
$\boldsymbol{f}=\boldsymbol{\Psi}\boldsymbol{F}\boldsymbol{D}_{\hat{\psi}}\boldsymbol{\hat{f}}.$ (4.9)
@@ 5 Comparison to the classical NFFT
$f(\boldsymbol{x})=\int_{\mathbb{R}^{d}}\hat{f}(\boldsymbol{v})\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{v}\boldsymbol{x}}\,\mathrm{d}\boldsymbol{v}=\int\limits_{[-\frac{M}{2},\frac{M}{2}]^{d}}\hat{f}(\boldsymbol{v})\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{v}\boldsymbol{x}}\,\mathrm{d}\boldsymbol{v}\approx\sum_{\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}}\hat{f}(\boldsymbol{k})\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{x}},$
such that the function evaluations $f(\boldsymbol{x}_{j})$, $j=1,\dots,N$, could also be approximated efficiently by means of an NFFT, see Algorithm 2.1. Since this raises the question of which of the two methods, Algorithm 2.1 or Algorithm 4.1, is more advantageous, this section deals with the comparison of the two approaches.
Considering the matrix notations $\boldsymbol{B}\boldsymbol{F}\boldsymbol{D}$ and $\boldsymbol{\Psi}\boldsymbol{F}\boldsymbol{D}_{\hat{\psi}}$, cf. (2.6) and (4.9), the first thing to realize is that for $\boldsymbol{B}\in\mathbb{R}^{N\times|\mathcal{I}_{\boldsymbol{L}}|}$ in (2.4) the window function $\varphi_{m}(\boldsymbol{x})$ is used, while for $\boldsymbol{\Psi}\in\mathbb{R}^{N\times|\mathcal{I}_{\boldsymbol{L}}|}$ in (4.8) we consider the $\mathrm{sinc}$ regularized window function $\psi(\boldsymbol{x})$. A similar remark can also be made about the diagonal matrices $\boldsymbol{D}\in\mathbb{C}^{|\mathcal{I}_{\boldsymbol{M}}|\times|\mathcal{I}_{\boldsymbol{M}}|}$ in (2.2) and $\boldsymbol{D}_{\hat{\psi}}\in\mathbb{C}^{|\mathcal{I}_{\boldsymbol{M}}|\times|\mathcal{I}_{\boldsymbol{M}}|}$ in (4.7).
Additionally, it is important to note that the two methods can only be compared for $\boldsymbol{x}\in[-\tfrac{1}{2}+\tfrac{m}{L},\tfrac{1}{2}-\tfrac{m}{L})^{d}$, as the approximation by Algorithm 4.1 is only reasonable in this case. This implies that the matrix $\boldsymbol{B}$ in (2.4) is, unlike usual, non-periodic in this setting, whereas the matrix $\boldsymbol{\Psi}$ in (4.8) is inherently non-periodic by definition.
@@
for $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$ fixed, cf. (2.6) with $L=M_{\sigma}$, where $\tilde{\varphi}_{m}(\boldsymbol{x})=\sum_{\boldsymbol{r}\in\mathbb{Z}^{d}}\varphi_{m}(\boldsymbol{x}+\boldsymbol{r})$ denotes the $1$-periodic version of the compactly supported window function $\varphi_{m}$. Thus, we look for a comparable approximation of the exponential function using our newly proposed method in Algorithm 4.1. For this purpose, note that $g(\boldsymbol{x})\coloneqq\hat{\psi}(\boldsymbol{x})\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{x}}$ with $\boldsymbol{k}\in\mathbb{R}^{d}$ fixed, possesses the Fourier transform $\hat{g}(\boldsymbol{v})=\psi(\boldsymbol{k}-\boldsymbol{v})$. Therefore, we have $g\in\mathcal{B}_{M/2}(\mathbb{R}^{d})$ for all $\boldsymbol{k}\in\big{[}-\frac{M}{2}+\frac{m}{L},\frac{M}{2}-\frac{m}{L}\big{]}^{d}$, i. e., considering (3.5) for this function $g$ yields
$g(\boldsymbol{x})=\hat{\psi}(\boldsymbol{x})\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{x}}\approx\sum_{{\boldsymbol{\ell}}\in\mathbb{Z}^{d}}\hat{\psi}\big{(}\tfrac{{\boldsymbol{\ell}}}{L}\big{)}\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{\ell}/L}\,\psi\big{(}{\boldsymbol{x}}-\tfrac{{\boldsymbol{\ell}}}{L}\big{)},\quad\boldsymbol{x}\in\mathbb{R}^{d},$
@@
For the sake of simplicity we restrict ourselves to the one-dimensional setting $d=1$ for the visualization. To estimate the quality of the approaches, we consider the approximation error
@@
where the term $h(x_{p})$ is a placeholder for the right hand sides of (5.1) and (5.2), respectively, evaluated at a fine grid of $P=10^{5}$ equispaced points $x_{p}$, $p=1,\dots,P$. This approximation error (5.3) shall now be computed for several values
$v_{s}=-\tfrac{M}{2}-m+\tfrac{s}{S}\in\big{[}-\tfrac{M}{2}-m,\tfrac{M}{2}+m\big{]},\quad s=0,\dots,S(M+2m),$ (5.4)
where $S=1$ corresponds to integer evaluation, whereas we use $S=32$ to examine the approximation at non-integer points as well. Note that (5.1) is expected to provide a good approximation only for $v\in\big{[}-\frac{M}{2},\frac{M}{2}\big{]}$, while (5.2) is expected to do so only for $v\in\big{[}-\frac{M}{2}+\frac{m}{L},\frac{M}{2}-\frac{m}{L}\big{]}$. Nevertheless, we test for $v$ from a larger interval to confirm this assumption.
@@
That is to say, Figure 5.1 demonstrates that the novel NFFT-like approach in Algorithm 4.1 is better suited for bandlimited functions, while this superiority is not limited to $k\in\mathcal{I}_{M}$ but extends to the entire domain $v\in\big{[}-\frac{M}{2},\frac{M}{2}\big{]}$. Moreover, the error of Algorithm 4.1 is bounded by the error estimates of the regularized Shannon sampling formulas in Section 3, whereas the quadrature error of the NFFT is completely unclear.
@@
Figure 5.1: Maximum approximation error (5.3) for $P=10^{5}$ computed for (5.4) with $S=32$ using the $\sinh$-type window function (3.4) as well as the parameters $M=20$, $\lambda=1$, $L=(1+\lambda)M$, and $m=5$ in the one-dimensional setting $d=1$.
@@
To examine the approximation quality of the NFFT-like procedure for bandlimited functions in Algorithm 4.1 we provide a function $f$ with its corresponding Fourier transform $\hat{f}$ in (3.1), such that we have access to the exact values $\hat{f}(k)$, $k\in\mathcal{I}_{M}$, as input for Algorithm 4.1, as well as the exact function evaluations $f(x_{j})$, $j=1,\dots,N$. In doing so, we can compare the result $f_{j}$, $j=1,\dots,N$, of Algorithm 4.1 to the exact function evaluations $f(x_{j})$, $j=1,\dots,N$, by computing the maximum approximation error
@@
For the reason of comparison, we also compute the approximation error (5.5) when $f_{j}$, $j=1,\dots,N$, is the result of the classical NFFT in Algorithm 2.1.
Here we consider the one-dimensional setting with $d=1$ and for several bandwidth parameters $M\in\{20,40,\dots,1000\}$ we study the function $f(x)=\mathrm{sinc}^{2}\big{(}\frac{M}{2}\pi x\big{)}$ with the Fourier transform
@@
Note that the function $f$ is scaled appropriately such that $\max_{x\in\mathbb{R}}f(x)=1$ independent of the bandwidth $M$ and thereby the approximation errors (5.5) are comparable for all considered $M$. As evaluation points $x_{j}\in\big{[}-\frac{1}{2}+\frac{m}{L},\,\frac{1}{2}-\frac{m}{L}\big{]}$, $j=1,\dots,N$, we choose the scaled Chebyshev nodes
$x_{j}=\cos\left(\frac{(j-1)\pi}{N}\right)\cdot\bigg{(}\frac{1}{2}-\frac{m}{L}\bigg{)},\quad j=1,\dots,N,$ (5.6)
@@
The corresponding results are depicted in Figure 5.2. As expected by the previous comparison of the two approaches in Figure 5.1, the new NFFT-like procedure for bandlimited functions presented in Algorithm 4.1 performs much better than the classical NFFT in Algorithm 2.1. While for $M\leq 80$ both approaches exhibit the same maximum approximation error (5.5), for larger bandwidth $M$ the approximation error (5.5) gets smaller only for the NFFT-like procedure in Algorithm 4.1. That is to say, when approximating the evaluations $f(x_{j})$, $j=1,\dots,N$, of the bandlimited function $f\in\mathcal{B}_{M/2}(\mathbb{R})$ by given samples $\hat{f}(k)$, $k\in\mathcal{I}_{M}$, of the corresponding Fourier transform (3.1), reasonable results can be obtained by the NFFT in Algorithm 2.1, yet evidence indicates that our newly proposed NFFT-like procedure for bandlimited functions in Algorithm 4.1 yields results that are at least as good, if not superior. Accordingly, the NFFT-like procedure in Algorithm 4.1 is the preferred approach in this context.
Figure 5.2: Maximum approximation error (5.5) of the NFFT-like procedure for bandlimited functions in Algorithm 4.1 and the classical NFFT in Algorithm 2.1 using the $\sinh$-type window function (3.4) computed for the function $f(x)=\mathrm{sinc}^{2}\big{(}\frac{M}{2}\pi x\big{)}$ using several bandwidth parameters $M\in\{20,40,\dots,1000\}$ and the scaled Chebyshev nodes (5.6) with $N=\frac{M}{2}$, $m=5$, $M_{\sigma}=L=(1+\lambda)M$, as well as $\lambda=1$ and $d=1$.

### v2

@@ 4 NFFT-like procedure for bandlimited functions
$\boldsymbol{D}_{\hat{\psi}}\coloneqq\text{diag}\left(\frac{1}{|\mathcal{I}_{\boldsymbol{L}}|\cdot\hat{\psi}(\boldsymbol{k})}\right)_{\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}}\ \in\mathbb{C}^{|\mathcal{I}_{\boldsymbol{M}}|\times|\mathcal{I}_{\boldsymbol{M}}|}$ (4.9)
@@
$\boldsymbol{f}=\boldsymbol{\Psi}\boldsymbol{F}\boldsymbol{D}_{\hat{\psi}}\boldsymbol{\hat{f}},$ (4.11)
@@ 5 Comparison to the classical NFFT
$f(\boldsymbol{x})=\int\limits_{[-\frac{M}{2},\frac{M}{2}]^{d}}\hat{f}(\boldsymbol{v})\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{v}\boldsymbol{x}}\,\mathrm{d}\boldsymbol{v}\approx\sum_{\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}}\hat{f}(\boldsymbol{k})\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{x}},$
such that the function evaluations $f(\boldsymbol{x}_{j})$, $j=1,\dots,N$, could also be approximated efficiently by means of an NFFT. Since this raises the question of which of the two methods, Algorithm 2.1 or Algorithm 4.2, is more advantageous, this section deals with the comparison of the two approaches.
Considering the matrix notations $\boldsymbol{B}\boldsymbol{F}\boldsymbol{D}$ and $\boldsymbol{\Psi}\boldsymbol{F}\boldsymbol{D}_{\hat{\psi}}$, cf. (2.6) and (4.11), the first thing to realize is that for $\boldsymbol{B}\in\mathbb{R}^{N\times|\mathcal{I}_{\boldsymbol{L}}|}$ in (2.4) the window function $\varphi_{m}(\boldsymbol{x})$ is used, while for $\boldsymbol{\Psi}\in\mathbb{R}^{N\times|\mathcal{I}_{\boldsymbol{L}}|}$ in (4.10) we consider the regularized $\mathrm{sinc}$ function $\psi(\boldsymbol{x})$ in (4.2). A similar remark can also be made about the diagonal matrices $\boldsymbol{D}\in\mathbb{C}^{|\mathcal{I}_{\boldsymbol{M}}|\times|\mathcal{I}_{\boldsymbol{M}}|}$ in (2.2) and $\boldsymbol{D}_{\hat{\psi}}\in\mathbb{C}^{|\mathcal{I}_{\boldsymbol{M}}|\times|\mathcal{I}_{\boldsymbol{M}}|}$ in (4.9).
Additionally, it is important to note that the two methods can only be compared for $\boldsymbol{x}\in[-\tfrac{1}{2}+\tfrac{m}{L},\tfrac{1}{2}-\tfrac{m}{L})^{d}$, as the approximation by Algorithm 4.2 is only reasonable in this case. This implies that the matrix $\boldsymbol{B}$ in (2.4) is, unlike usual, non-periodic, whereas the matrix $\boldsymbol{\Psi}$ in (4.10) is inherently non-periodic by definition.
@@
for $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{M}}$ fixed, cf. (2.6) with $L=M_{\sigma}$, where $\tilde{\varphi}_{m}(\boldsymbol{x})=\sum_{\boldsymbol{r}\in\mathbb{Z}^{d}}\varphi_{m}(\boldsymbol{x}+\boldsymbol{r})$ denotes the $1$-periodic version of the compactly supported window function $\varphi_{m}$. Thus, we look for a comparable approximation of the exponential function using our newly proposed method in Algorithm 4.2. For this purpose, note that $g(\boldsymbol{x})\coloneqq\hat{\psi}(\boldsymbol{x})\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{x}}$ with $\boldsymbol{k}\in\mathbb{R}^{d}$ fixed possesses the Fourier transform $\hat{g}(\boldsymbol{v})=\psi(\boldsymbol{k}-\boldsymbol{v})$. Therefore, we have $g\in\mathcal{B}_{M/2}(\mathbb{R}^{d})$ for all $\boldsymbol{k}\in\big{[}-\frac{M}{2}+\frac{m}{L},\frac{M}{2}-\frac{m}{L}\big{]}^{d}$, i. e., considering (3.5) for this function $g$ yields
$\hat{\psi}(\boldsymbol{x})\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{x}}\approx\sum_{{\boldsymbol{\ell}}\in\mathbb{Z}^{d}}\hat{\psi}\big{(}\tfrac{{\boldsymbol{\ell}}}{L}\big{)}\,\mathrm{e}^{2\pi\mathrm{i}\boldsymbol{k}\boldsymbol{\ell}/L}\,\psi\big{(}{\boldsymbol{x}}-\tfrac{{\boldsymbol{\ell}}}{L}\big{)},\quad\boldsymbol{x}\in\mathbb{R}^{d},$
@@
For simplicity we restrict ourselves to the one-dimensional setting $d=1$ for the visualization. To estimate the quality of the approaches, we consider the approximation error
@@
where the term $h(x_{p})$ is a placeholder for the right-hand sides of (5.1) and (5.2), respectively, evaluated at a fine grid of $P=10^{5}$ equispaced points $x_{p}$, $p=1,\dots,P$. This approximation error (5.3) shall now be computed for several values
$\hskip 20.00003ptv_{s}$ $=-\tfrac{M}{2}-m+\tfrac{s}{S}\in\big{[}-\tfrac{M}{2}-m,\tfrac{M}{2}+m\big{]},\quad s=0,\dots,S(M+2m),$ (5.4)
where $S=1$ corresponds to integer evaluation, whereas we use $S=32$ to examine the approximation at non-integer points as well. Note that (5.1) is expected to provide a good approximation only for $v\in\big{[}-\frac{M}{2},\frac{M}{2}\big{]}$, while (5.2) is expected to do so only for $v\in\big{[}-\frac{M}{2}+\frac{m}{L},\frac{M}{2}-\frac{m}{L}\big{]}$. Nevertheless, we test for $v$ from a larger interval to confirm these assumptions.
@@
That is to say, Figure 5.1 demonstrates that the novel NFFT-like approach in Algorithm 4.2 is better suited for bandlimited functions, while this superiority is not limited to $k\in\mathcal{I}_{M}$ but extends to the entire domain $v\in\big{[}-\frac{M}{2},\frac{M}{2}\big{]}$. Moreover, the error of Algorithm 4.2 is bounded by the error estimates of the regularized Shannon sampling formulas in Section 3, whereas the quadrature error of the NFFT remains unclear.
@@
Figure 5.1: Maximum approximation error (5.3) for $P=10^{5}$ computed for (5.4) with $S=32$ using the $\sinh$-type window function (3.4) as well as $M=20$, $\lambda=1$, $L=(1+\lambda)M$, and $m=5$ in the one-dimensional setting $d=1$.
@@
To finally examine the approximation quality of the NFFT-like procedure in Algorithm 4.2 for bandlimited functions we provide a function $f$ with its corresponding Fourier transform $\hat{f}$ in (3.1), such that we have access to the exact values $\hat{f}(k)$, $k\in\mathcal{I}_{M}$, as input for Algorithm 4.2, as well as the exact function evaluations $f(x_{j})$, $j=1,\dots,N$. In doing so, we can compare the result $f_{j}$, $j=1,\dots,N$, of Algorithm 4.2 to the exact function evaluations $f(x_{j})$, $j=1,\dots,N$, by computing the maximum approximation error
@@
For comparison we also compute the approximation error (5.5) when $f_{j}$ is the result of the classical NFFT in Algorithm 2.1.
We consider the one-dimensional setting $d=1$ and for several bandwidth parameters $M\in\{20,40,\dots,1000\}$ we study the function $f(x)=\mathrm{sinc}^{2}\big{(}\frac{M}{2}\pi x\big{)}$ with the Fourier transform
@@
Note that the function $f$ is scaled such that $\max_{x\in\mathbb{R}}f(x)=1$ independent of the bandwidth $M$ and thereby the approximation errors (5.5) are comparable for all considered $M$. As evaluation points $x_{j}\in\big{[}-\frac{1}{2}+\frac{m}{L},\,\frac{1}{2}-\frac{m}{L}\big{]}$, $j=1,\dots,N$, we choose the scaled Chebyshev nodes
$x_{j}=\cos\left(\frac{(j-1)\pi}{N}\right)\cdot\left(\frac{1}{2}-\frac{m}{L}\right),\quad j=1,\dots,N,$ (5.6)
@@
The corresponding results are depicted in Figure 5.2. As expected by Figure 5.1, the new NFFT-like procedure in Algorithm 4.2 performs much better than the classical NFFT in Algorithm 2.1. While for $M\leq 80$ both approaches exhibit the same maximum approximation error (5.5), for larger bandwidth $M$ the approximation error (5.5) gets smaller only for the NFFT-like procedure in Algorithm 4.2. That is to say, when approximating the evaluations $f(x_{j})$, $j=1,\dots,N$, of the bandlimited function $f\in\mathcal{B}_{M/2}(\mathbb{R})$ by given samples $\hat{f}(k)$, $k\in\mathcal{I}_{M}$, of the corresponding Fourier transform (3.1), reasonable results can be obtained by the NFFT in Algorithm 2.1, yet evidence indicates that our newly proposed NFFT-like procedure for bandlimited functions in Algorithm 4.2 yields results that are at least as good, if not superior. Accordingly, we conclude that the NFFT-like procedure in Algorithm 4.2 is the preferred approach in this context.
Figure 5.2: Maximum approximation error (5.5) of Algorithms 2.1 and 4.2 using the $\sinh$-type window function (3.4) computed for the function $f(x)=\mathrm{sinc}^{2}\big{(}\frac{M}{2}\pi x\big{)}$, $M\in\{20,40,\dots,1000\}$, and the scaled Chebyshev nodes (5.6) with $N=\frac{M}{2}$, $m=5$, $M_{\sigma}=L=(1+\lambda)M$, as well as $\lambda=1$ and $d=1$.
Acknowledgments
[…]